OpenStudy (anonymous):

what is velocity of ring frolling form inclined plane without slipping ?give expression? plzz

OpenStudy (anonymous):

Use the conservation of energy.Suppose ring rolls down a height of h and suppose it had zero velocity initially. Then mgh = K.E. Remember that K.E. is not the kinetic energy only due to translation(1/2mv^2) but also kinetic energy due to rotation(1/2* I *w^2) w= angular velocity as the object is rolling w = v/r where r is radius of ring. I =mr^2